# Multi-Modal and Multi-Agent Systems Meet Rationality: A Survey

### v2

@@ 6 Open Problems and Future Directions
It is important to understand that integrating most of these agents or modules with LLMs still does not inherently make LLMs more rational. Current methods are neither sufficient nor necessary, but they serves as instrumental tools that bridges the gap between an LLM’s response and rationality. These approaches enables multi-agent systems, which are black boxes from the user’s perspective, to more closely mimic rational thinking in their output responses. However, despite these more rational responses elicited from multi-modal and multi-agent systems, the challenge of how to effectively close the loop and bake these enhanced outputs back into the LLMs Zhao et al. (2024), beyond mere fine-tuning, remains an open topic. In other words, can we leverage these more rational outputs to inherently enhance a single foundation model’s rationality in its initial responses in future applications?
@@
Future research should prioritize the development of more robust and scalable methods for evaluating rationality, taking into account unique challenges and biases posed by agents. A promising direction is to create benchmarks specifically tailored to assess rationality, going beyond existing ones on accuracy. These new benchmarks should avoid data contamination and emphasize tasks that demand consistent reasoning across diverse representations and domains. There is a need for more rigorous and large-scale studies on the principles of invariance and orderability of preference, together with their applications to testing rationality in agent systems. This would involve developing more sophisticated perturbation methods that probe the consistency of reasoning at a deeper level, as well as designing experiments that yield statistically significant results.
@@ 7 Conclusions
This survey builds connections between multi-modal and multi-agent systems with rationality, guided by dual-process theories and the four axioms we expect a rational agent or agent systems should satisfy: information grounding, orderability of preference, independence from irrelevant context, and invariance across logically equivalent representations. Our findings suggest that the grounding can usually be enhanced by multi-modalities, world models, knowledge retrieval, and tool utilization. The remaining three axioms are typically intertwined, and we sometimes describe their collective characteristics informally using terms such as coherence, consistency, and trustworthiness. These axioms are simultaneously improved by achievements in multi-modalities, tool utilization, neuro-symbolic reasoning, self-reflection, and multi-agent collaborations. These fields of research, by either slowing down the “thinking" process or boiling down real-world problems to logical essence, mimic the “System 2" thinking in human cognition, thereby enhancing the rationality of multi-agent systems in decision-making scenarios, compared to single-agent language-only baselines that resemble the “System 1" process.
@@ 8 Limitations
Appendix B The Orderability of Preferences Matters for LLM-based Evaluations
@@
This section talks about LLM-based evaluation rather than evaluating the rationality of LLMs discussed in Section 5. Recent research underscores a critical need for more rational LLM-based evaluation methods, particularly for assessing open-ended language responses. CoBBLEr Koo et al. (2023) provides a cognitive bias benchmark for evaluating LLMs as evaluators, revealing a preference for their own outputs over those from other LLMs. Stureborg et al. (2024) argues that LLMs are biased evaluators towards more familiar tokens and previous predictions, and exhibit strong self-inconsistency in the score distribution. Luo et al. (2023); Shen et al. (2023); Gao et al. (2023c); Wang et al. (2023b); Chen et al. (2023); Chiang and Lee (2023); Zheng et al. (2024b); Fu et al. (2023); Liu et al. (2023c) also point out the problem with a single LLM as the evaluator, with concerns over factual and rating inconsistencies, a high dependency on prompt design, a low correlation with human evaluations, and struggles with the comparison. As a result, having a coherent orderability of preferences aligned with human preference becomes increasingly important.

### v3

@@ 6 Open Problems and Future Directions
It is important to understand that the notion of multi-modal or multi-agent systems does not inherently imply rationality. Current methods are neither sufficient nor necessary, but they serve as instrumental tools that bridge the gap between an LLM’s response and rationality. These approaches enable multi-agent systems, which are black boxes from the user’s perspective, to more closely mimic rational thinking in their output responses. However, despite these more rational responses elicited from multi-modal and multi-agent systems, the challenge of how to effectively close the loop and bake these enhanced outputs back into foundation models themselves Zhao et al. (2024) beyond mere fine-tuning remains an open question. In other words, can we leverage these more rational outputs to inherently enhance a single foundation model’s rationality in its initial responses in future applications?
@@
Future research should prioritize the development of more robust and scalable methods for evaluating rationality, taking into account unique challenges and biases Jiang et al. (2024a) posed by agents. A promising direction is to create methods specifically tailored to assess rationality, going beyond existing ones on accuracy. These new methods should avoid data contamination and emphasize tasks that demand consistent reasoning across diverse representations and domains. There is a need for more rigorous and large-scale studies on the principles of invariance and orderability of preference, together with their applications to testing rationality in agent systems. This would involve developing more sophisticated perturbation methods that probe the consistency of reasoning at a deeper level, as well as designing experiments that yield statistically significant results.
@@ 7 Conclusions
This survey builds connections between multi-modal and multi-agent systems with rationality, guided by dual-process theories and the four axioms we expect a rational agent or agent systems should satisfy: grounding, orderability of preference, independence from irrelevant context, and invariance. Our findings suggest that the grounding can usually be enhanced by multi-modalities, knowledge retrieval, and tool utilization. The remaining three axioms are typically intertwined, and often simultaneously improved via deliberation (slow, iterative thinking process) and abstraction (distilling the logical essence).
@@ 8 Limitations
Appendix B Information Grounding
Web agents are a quintessential example of how multi-modal agents surpass language-only ones. In agents like Pix2Act Shaw et al. (2024), WebGUM Furuta et al. (2023), CogAgentHong et al. (2023b), and SeeAct Zheng et al. (2024a), web navigation is grounded on graphical user interface (GUI) rather than solely on HTML texts Shen et al. (2024a); Yao et al. (2022a); Deng et al. (2024); Gur et al. (2023). This method of visual grounding offers higher information density compared to HTML codes that are usually lengthy, noisy, and sometimes even incomplete Zheng et al. (2024a). Supporting the importance of vision, ablation studies in WebGUM Furuta et al. (2023) also reports $5.5\%$ success rate improvement on the MiniWoB++ dataset Shi et al. (2017); Liu et al. (2018) by simply adding the image modality.
Multi-modalities also help enhance the functionality of agent systems through more diverse information grounding. For example, Chain-of-Action Pan et al. (2024) advances the single-modal Search-in-the-Chain Xu et al. (2023) by supporting multi-modal data retrieval for faithful question answering. DoraemonGPT Yang et al. (2024) decomposes complex tasks into simpler ones toward understanding dynamic scenes, where multi-modal understanding is necessary for spatial-temporal videos analysis. RA-CM3 Yasunaga et al. (2022) augments baseline retrieval-augmented LLMs with raw multi-modal documents that include both images and texts, assuming that these two modalities can contextualize each other and make the documents more informative, leading to better generator performance. The multi-modal capabilities also allow HuggingGPT Shen et al. (2024b), Agent LUMOS Yin et al. (2023), ToolAlpaca Tang et al. (2023), and AssistGPT Gao et al. (2023b) to expand the scope of tasks they can address, including cooperation among specialized agents or tools capable of handling different information modalities.
Large world models is an emerging and promising direction to reduce multi-modal hallucinations. The notion is also mentioned in “Objective-driven AI" LeCun (2024), where agents have behavior driven by fulfilling objectives and they understand how the world works with common sense knowledge, beyond an auto-regressive generation. For example, Large World Model (LWM) Liu et al. (2024b) and Sora Brooks et al. (2024) develop insights from both textual knowledge and the world through video sequences. Although these models both advance toward general-purpose simulators of the world, they still lack reliable physical engines for guaranteed grounding in real-world dynamics. Ghost-in-the-Minecraft Zhu et al. (2023b) and Voyager Wang et al. (2023a) have agents living in a well-defined game-world environment. JEPA LeCun (2022) creates a recurrent world model in an abstract representation space.
Appendix C Knowledge Retrieval & Tool Usage
CuriousLLM Yang and Zhu (2024) presents ablation studies showing the effectiveness of KGs on improving reasoning within the search process. MineDojo Fan et al. (2022) observes that internet-scale multi-modal knowledge allows models to significantly outperform all creative task baselines. Equipped with world knowledge, RA-CM3 Yasunaga et al. (2022) can finally generate faithful images from captions compared to CM3 Aghajanyan et al. (2022) and Stable Diffusion Rombach et al. (2022). CooperKGC Ye et al. (2023) enables multi-agent collaborations, leveraging knowledge bases of different experts. It finds that the incorporation of KGs improves F1 scores by $10.0$-$33.6\%$ across different backgrounds, and adding more collaboration rounds also enhance performance by about $10.0$-$30.0\%$. DoraemonGPT Yang et al. (2024) supports knowledge tools to assist the understanding of specialized video contents. SIRI Wang et al. (2023d) builds a multi-view knowledge base to increase the explainability of visual question answering.
A multi-agent system can coordinate agents understanding when and which tool to use, which modality of information the tool should expect, how to call the corresponding API, and how to incorporate outputs from the API calls, which anchors subsequent reasoning processes with more accurate information beyond their parametric memory. For example, VisProg Gupta and Kembhavi (2023), ViperGPT Surís et al. (2023), and Parsel Zelikman et al. (2023) generate Python programs to reliably execute subroutines. Gupta and Kembhavi (2023); Surís et al. (2023) also invoke off-the-shelf models for multimodal assistance.
Foundation models are not specifically trained for object detection or segmentation, so BuboGPT Zhao et al. (2023) and Multi-Agent VQA Jiang et al. (2024b) call SAM Kirillov et al. (2023); Ren et al. (2024) as the tool. Besides, BabyAGI Nakajima (2023), Chamelon Lu et al. (2024), AssistGPT Gao et al. (2023b), Avis Hu et al. (2024), ToolAlpaca Tang et al. (2023), MetaGPT Hong et al. (2023a), Agent LUMOS Yin et al. (2023), AutoAct Qiao et al. (2024), $\alpha$-UMi Shen et al. (2024a), and ConAgents Shi et al. (2024) harness compositional reasoning to enable generalized multi-agent systems with planning and modular tool-using capabilities in real-world scenarios.
In most cases, tools require translating natural language queries into API calls with predefined syntax. Once the APIs and their input arguments are determined, the tools will ignore any irrelevant context in the original queries, as long as the queries share the same underlying logic necessary for the inputs. Take Multi-Agent VQA Jiang et al. (2024b) as an example. In this system, a language model provides only the relevant object names to the Grounded SAM Ren et al. (2024) component, which functions as an object detector, rather than passing the entire visual question. Similarly, the image editing tools in VisProg Gupta and Kembhavi (2023) only receive a fixed set of arguments translated from user queries to perform deterministic code executions. SeeAct Zheng et al. (2024a) as a Web agent explores vision-language models, ranking models, and a bounding box annotation tool to improve Web elements grounding from lengthy and noisy HTML codes.
Appendix D Collective Deliberation among Agents
D.1 More Examples on Multi-Agent Collaborations
Corex Sun et al. (2023a) finds that orchestrating multiple agents to work together yields better complex reasoning results, exceeding strong single-agent baselines Wang et al. (2022b) by an average of $1.1$-$10.6\%$. Retroformer Yao et al. (2023) equips the single-agent Reflexion Shinn et al. (2024) algorithm with an additional LLM to generate verbal reinforcement cues and assist its self-improvement, enhancing accuracy by $1.0$-$20.9\%$. MetaAgents Li et al. (2023i) effectively coordinate agents within task-oriented social contexts to achieve consistent behavior patterns, and the implementation of agent reflection in this system leads to a $21.0\%$ improvement in success rates. Multi-agent debating in Khan et al. (2024) also leads to more truthful answers, boosting single-agent baselines by $28.0\%$. Multi-Agent Collaboration Talebirad and Nadiri (2023), ChatDev Qian et al. (2023), AgentCF Zhang et al. (2023), AutoGen Wu et al. (2023), Social Learning Mohtashami et al. (2023), S3 Gao et al. (2023a), Ke et al. (2024), and Chern et al. (2024) continue to push the frontier of a multi-agent system’s applications beyond daily conversation to a versatile set of real-world task completions.
D.2 Collaboration Againt Jailbreaking
LLMs are also sensitive to prompt perturbations due to token bias and noises Sclar et al. (2023a). One of the most worrying examples are adversarial attacks Gehman et al. (2020); Ganguli et al. (2022); Du et al. (2022); Wei et al. (2024); Perez et al. (2022); Zou et al. (2023) through malicious prompt engineering. These attacks, also known as the Red Team Task, also named the Red Team Task, involve malicious prompt engineering designed to exploit vulnerabilities in the model. To combat this issue, Chern et al. (2024) propose a multi-agent debating approach involving agents with harmless, neutral, or harmful intentions. The authors demonstrate that engaging these agents in multi-round, multi-agent debate is more effective in improving the model’s robustness against adversarial prompt variations and perturbations compared to a single-agent with self-reflection prompts.
D.3 Collaboration on LLM-based Evaluation
LLM-based evaluation methods are popular in assessing open-ended language responses. Stureborg et al. (2024); Koo et al. (2023) point out LLMs often present cognitive biases in their evaluations, favoring certain types of responses over others regardless of the actual quality or relevance of the respective responses. To establish a more coherent preference orderability aligned with human preference. ChatEval Chan et al. (2023) introduces a multi-agent debate framework to mimic human annotators collaborating in robust answer evaluations. Its multi-agent approach achieves greater alignment with human preferences compared to single-agent evaluations, enhancing accuracy by $6.2\%$ for GPT-3.5 and $2.5\%$ for GPT-4, and an increase of $16.3\%$ and $10.0\%$ in average Spearman and Kendall-Tau correlations Zhong et al. (2022) with human judgements in GPT-4.
D.4 The Orderability of Preferences Matters for LLM-based Evaluations
@@
This section talks about LLM-based evaluation rather than evaluating the rationality of LLMs discussed in Section 5. Recent research underscores a critical need for more rational LLM-based evaluation methods, particularly for assessing open-ended language responses. CoBBLEr Koo et al. (2023) provides a cognitive bias benchmark for evaluating LLMs as evaluators, revealing a preference for their own outputs over those from other LLMs. Stureborg et al. (2024) argues that LLMs are biased evaluators towards more familiar tokens and previous predictions, and exhibit strong self-inconsistency in the score distribution. Luo et al. (2023); Shen et al. (2023); Gao et al. (2023c); Wang et al. (2023b); Chen et al. (2023); Chiang and Lee (2023); Zheng et al. (2024b); Fu et al. (2023); Liu et al. (2023b) also point out the problem with a single LLM as the evaluator, with concerns over factual and rating inconsistencies, a high dependency on prompt design, a low correlation with human evaluations, and struggles with the comparison. As a result, having a coherent orderability of preferences aligned with human preference becomes increasingly important.
@@
Appendix E Neuro-Symbolic Reasoning
Logic-LM Pan et al. (2023) combines problem formulating, symbolic reasoning, and result interpreting agents, where the symbolic reasoner empowers LLMs with deterministic symbolic solvers to perform inference, ensuring a correct answer is consistently chosen. Its multi-agent framework also encourages self-refinement that modifies logical formulation errors using error messages from the symbolic reasoner as the feedback. Besides, SymbolicToM Sclar et al. (2023b) and KRISP Marino et al. (2021) construct explicit symbolic graphs and answer questions by retrieving nodes in the graph. Binder Cheng et al. (2022), Parsel Zelikman et al. (2023), LEFT Hsu et al. (2024), and Fang et al. (2024) decompose tasks into planning, parsing, and execution, where the symbolic reasoning agents can help maintain a coherent order of preferences among symbolic options in the system outputs.
Appendix F Evaluating Rationality
F.1 Benchmarks for Hallucination
Multiple evaluation benchmarks targeting language-only dialogue have been proposed, such as BEGIN Dziri et al. (2022b), HaluEval Li et al. (2023e), DialFact Gupta et al. (2021), FaithDial Dziri et al. (2022a), AIS Rashkin et al. (2023), and others Zheng et al. (2023b); Das et al. (2023); Cao et al. (2021). In contrast, benchmarks on multi-agent frameworks beyond language dialogue or those involving multi-modalities are very limited. Liu et al. (2024a) moves beyond conversation to code generation; EureQA Li et al. (2023a) focuses on reasoning chains; and TofuEval Tang et al. (2024) evaluates hallucination in multi-domain summarization. Object hallucination Rohrbach et al. (2018); Biten et al. (2022), POPE Li et al. (2023g), and LLaVA-RLHF Sun et al. (2023b) are the few examples evaluating multi-modal hallucination.
F.2 Perturbation Techniques
Perturbation techniques typically involve some versions of paraphrasing or permutation. Paraphrasing includes changing the instruction templates Weber et al. (2023), rewording task descriptions (Yang et al., 2023; Ohmer et al., 2024; Wang et al., 2024b), translating the prompts into a different language (Ohmer et al., 2023, 2024; Xu et al., 2024a) and then back to the original language Yang et al. (2023), and making subtle changes to entities in task descriptions without affecting the logical structure, like altering names of the characters, numerical values in math problems, or locations of the events Wang et al. (2024b). Permutation also includes reordering in-context learning examples Lu et al. (2021); Pecher et al. (2024) and, in the case of multiple-choice questions, rearranging the options (Zong et al., 2023; Zheng et al., 2023a).